# Improving exponential-family random graph models for bipartite networks

### v1

@@ 5 New statistics for modelling four-cycles in bipartite ERGMs
Let
$L_{2}(i,j)=\sum_{h\neq i,j}x_{ih}x_{hj}$ (2)
be the number of two-paths connecting node $i$ to node $j$. Since a four-cycle is a combination of two two-paths [30, p. 123], the number of four-cycles is
@@
where $d(i,j)$ is the geodesic distance from $i$ to $j$. Some illustrative examples of the value of $C_{4}(i)$ for different nodes in some small graphs are shown in Fig. 1.
@@
where $0<\alpha\leq 1$ is the exponent for, in the terminology of Wilson et al. [76], the “$\alpha$-inside” weighting, since the subgraph counts ($C_{4}(i)$ here) are exponentiated before summing over all subgraphs. The “$\alpha$-outside” weighting would be to exponentiate the statistic after summing over all subgraphs, that is, in this case it would be $\left[\sum_{i}C_{4}(i)\right]^{\alpha}$. As discussed in Wilson et al. [76], the $\alpha$-inside weighting leads to local dependence as usually used in ERGMs, while the $\alpha$-outside weighting leads to global dependence, in which all ties are dependent on each other to some degree [76, p. 41]. The nature of the local dependencies induced by the new change statistics defined here using the $\alpha$-inside weighting are discussed in Section 5.2 below.
@@
Figure 2: Representations of the new configurations (left to right) FourCyclesNodePower, BipartiteFourCyclesNodePowerA, and BipartiteFourCyclesNodePowerB. Nodes in node set A are shown as red circles, and nodes in node set B are shown as blue squares.
@@
The “four-fan-3” graph in Table 3 is the same graph as the representation of the BipartiteFourCyclesNodePowerA configuration in Fig. 2 (ignoring the vertical ellipsis so that exactly three four-cycles are present). Note, however, that this has the counterintuitive property that, although it is a representation of the BipartiteFourCyclesNodePowerA configuration, in fact the value of the BipartiteFourCyclesNodePowerB statistic is greater than that of the BipartiteFourCyclesNodePowerA statistic for this graph. This is because the value of the statistic [(8) or (9)] is the sum over all nodes in the relevant node set ($A$ or $B$, respectively) of the four-cycle count at each node (4) raised to the power $\alpha$ (“$\alpha$-inside” weighting). Therefore, in this graph, the nodes in mode $B$ contribute more to the total as each one (of the six) is involved in exactly one four-cycle (and hence raising to the power of $\alpha$ still contributes one to the sum), while of the four nodes in mode $A$, three are involved in only one four-cycle, while the fourth is involved in three four-cycles and hence contributes only $3^{\alpha}\approx 1.73205$ (when $\alpha=0.5$). Generalising this four-fan-3 graph to four-fan-$k$ ($k\geq 1$, and if $k=1$ the graph is just a four-cycle) with the central high-degree node in node set $A$, we have:
@@
Because $0<\alpha\leq 1$, the BipartiteFourCyclesNodePowerA statistic (16) will always be less than (or equal to, if $\alpha=1$ or $k=1$) the BipartiteFourCyclesNodePowerB statistic (17) for this family of graphs. The consequences for the interpretation of the corresponding parameters are discussed in Section 6.1.
@@ 5.1 Change statistics for the new statistics
$=\sum_{k\in N(j)}L_{2}(i,k)$ (19)
where $N(i)$ denotes the neighbours of node $i$, that is, nodes $k\neq i$ such that $x_{ik}=1$, or, equivalently, $d(i,k)=1$. The change statistic for the FourCyclesNodePower statistic (7) is then:
@@ 5.2 Position of the new statistics in the dependence hierarchy
The configurations allowed in a model are determined by the assumptions as to which ties are allowed to depend on which other ties. Pattison & Snijders [22] (subsequently elucidated by Wang et al. [19] for bipartite networks, and more recently by Pattison et al. [77]) created a two-dimensional hierarchy of dependence assumptions, where the two dimensions are two facets of proximity: the form of the proximity condition, and the maximum distance between dependent ties. This two-dimensional hierarchy of dependence assumptions is illustrated in Fig. 3, showing a partial order structure, in which, if one dependence condition can be implied by another, it can be reached by a downwards path from the first to the second [19, p. 215].
Figure 3: Dependence hierarchy, adapted from Wang et al. [19, Fig. 3]. $\mathrm{SI}_{0}$ is not defined.
In order to describe the proximity conditions, it is useful to define some notation for neighbourhoods in a graph. In Section 5.1 we defined $N(i)$ as the neighbours of node $i$, that is, nodes $k\neq i$ such that $d(i,k)=1$. Following Pattison et al. [77], we now define $N_{q}(u)$, the $q$-neighbourhood of node $u$, as the set of nodes within geodesic distance $q$ of $u$, that is, nodes $v$ such that $d(u,v)\leq q$. Note that $N(u)$ as earlier defined in Section 5.1 is distinct from $N_{1}(u)$, the former being defined as nodes with a geodesic distance of exactly 1 from $u$ (and hence excluding $u$ itself), while the latter includes $u$ itself as $d(u,u)=0$. Further, we define $N_{q}(U)$, the $q$-neighbourhood of node set $U$, as $N_{q}(U)=\{v:v\in N_{q}(u)\text{ for some }u\in U\}$, that is, the set of nodes whose distance to some node in $U$ is no more than $q$.
@@
The dependence condition $\mathrm{I}_{1}$ is equivalent to the widely-used “social circuit“ dependence assumption [19], in which two ties are conditionally dependent if they would form a four-cycle if both present [78, 79, 30, 37], and $\mathrm{PI}_{1}$ allows the “alternating pendant-triangle” statistics recently described by Pattison et al. [77].
@@
Proposition 3(d) of Pattison et al. [77] states that “For $\mathrm{D}_{\mathrm{p}}$, each configuration is a subgraph in which every pair of edges lies on a path of length $\leq(p+2)$[77, p. 191]. So for a configuration to be in $\mathrm{D}_{1}$, every pair of edges must lie on a path of length three (or shorter). The FourCyclesNodePower configuration (Fig. 2) does not meet this requirement, since there are edges that do not lie on a path of length three or less: consider, for example, a pair of edges incident to the outermost node in the figure on two different four-cycles. These do not lie on a path of length three (but are on a path of length four, satisfying the requirement for $\mathrm{D}_{2}$ but not $\mathrm{D}_{1}$). Hence the FourCyclesNodePower configuration is not in dependence class $\mathrm{D}_{1}$.
Proposition 3(c) of Pattison et al. [77] states that “For $\mathrm{PI}_{\mathrm{p}}$, each configuration is a subgraph in which every pair of edges lies either on a cyclic walk of length $\leq(2p+2)$ or on a cyclic walk of length $\leq 2(p-r)+1$ with an additional path of length $\leq r+1$ attached to a node lying on the cyclic walk, for $10\leq r\leq p-$ 1[77, p. 191]. So, for $\mathrm{PI}_{2}$, each configuration is a subgraph in which every pair of edges is on a cyclic walk of length $\leq(2p+2)=6$, or on a cyclic walk of length $\leq 2(p-0)+1=5$ with an additional path of length $\leq 1$ attached to a node on the cyclic walk, or on a cyclic walk of length $\leq 2(p-1)+1=3$ with an additional path of length $\leq 2$ attached to a node on the cyclic walk. Again, we can see that the configuration for FourCyclesNodePower does not meet these conditions, considering a pair of maximally distant edges (those incident to the outermost nodes in two different four-cycles in Fig. 2). Such a pair of edges is neither on a six-cycle, and nor is it on a five-cycle with an additional path of length at most one or a three-cycle with an additional path of length at most two. Hence the FourCyclesNodePower configuration is not in dependence class $\mathrm{PI}_{2}$.
In the case of the BipartiteFourCyclesNodePower ($A$ and $B$) statistics for bipartite networks, the same reasoning applies (see Fig. 2).
@@ 5.3 Implementation
The new ERGM effects FourCyclesNodePower, BipartiteFourCyclesNodePowerA, and BipartiteFourCyclesNodePowerB are implemented in the EstimNetDirected [71] software, available from https://github.com/stivalaa/EstimNetDirected. BipartiteFourCyclesNodePowerA and BipartiteFourCyclesNodePowerB are also implemented, as b1np4c and b2np4c, as user-contrubuted statnet moodel terms [80, 81], available from https://github.com/stivalaa/ergm.terms.contrib. An example is described in Appendix C.
The b1nodematch and b2nodematch statistics [21] are also implemented in EstimNetDirected (as BipartiteNodematchAlphaA, BipartiteNodematchAlphaB, BipartiteNodematchBetaA, and BipartiteNodematchBetaB). These statistics are described in Section 7.2.
@@ 6 Simulation experiments
The results of these simulations are shown in Fig. 4. Using the BipartiteAltKCyclesA parameter (left column) results in phase transition or near-degeneracy behaviour, with the statistic showing a sudden sharp increase at a critical value of the parameter. At this critical value, the graph density (top left plot) also sharply increases, as does the number of four-cycles (bottom left plot), which, at parameter values less than the critical value, hardly increased at all. This behaviour is similar to that of the simple Markov (edge-triangle) model described in Koskinen & Daraganova [37], and is characteristic of near-degeneracy in ERGMs. This can prevent estimation of models which contain parameters that cause this behaviour, and yet occurs in this case even when using an “alternating” statistic, designed to try to avoid such behaviour [18]. Note that changing the $\lambda$ parameter appears merely to change the maximum value of the statistic; it does not remove or “smooth out” the phase transition. In contrast, when using the new BipartiteFourCyclesNodePowerA parameter (right column) the phase transition behaviour is less apparent even for the highest value of $\alpha$ ($1/2$), and can be smoothed out further as we decrease $\alpha$. So, by appropriately setting $\alpha$, we can use the BipartiteFourCyclesNodePowerA parameter to generate smoothly varying numbers of four-cycles, without suddenly tipping from a low-density low-clustering regime to a high-density high-clustering regime with nothing in between, which did not seem to be possible on this example with the BipartiteAltKCyclesA parameter.

### v3

@@ 6 New statistics for modelling four-cycles in bipartite ERGMs
Since a four-cycle is a combination of two two-paths [33, p. 123], the number of four-cycles is
@@
Some illustrative examples of the value of $C_{4}(i)$ for different nodes in some small graphs are shown in Fig. 2.
@@
where $0<\alpha\leq 1$ is the exponent for, in the terminology of Wilson et al. [79], the “$\alpha$-inside” weighting, since the subgraph counts ($C_{4}(i)$ here) are exponentiated before summing over all subgraphs. The “$\alpha$-outside” weighting would be to exponentiate the statistic after summing over all subgraphs, that is, in this case it would be $\left[\sum_{i}C_{4}(i)\right]^{\alpha}$. As discussed in Wilson et al. [79], the $\alpha$-inside weighting leads to local dependence as usually used in ERGMs, while the $\alpha$-outside weighting leads to global dependence, in which all ties are dependent on each other to some degree [79, p. 41]. The nature of the local dependencies induced by the new change statistics defined here using the $\alpha$-inside weighting are discussed in Section 6.3 below.
@@
Figure 3: Representations of the new configurations (a) FourCyclesNodePower, (b) BipartiteFourCyclesNodePowerA, and (c) BipartiteFourCyclesNodePowerB. Nodes in node set A are shown as red circles, and nodes in node set B are shown as blue squares.
@@
The “four-fan-3” graph in Table 3 is the same graph as the representation of the BipartiteFourCyclesNodePowerA configuration in Fig. 3(b) (ignoring the vertical ellipsis so that exactly three four-cycles are present). Note, however, that this has the counterintuitive property that, although it is a representation of the BipartiteFourCyclesNodePowerA configuration, in fact the value of the BipartiteFourCyclesNodePowerB statistic is greater than that of the BipartiteFourCyclesNodePowerA statistic for this graph. This is because the value of the statistic [(14) or (15)] is the sum over all nodes in the relevant node set ($A$ or $B$, respectively) of the four-cycle count at each node (10) raised to the power $\alpha$ (“$\alpha$-inside” weighting). Therefore, in this graph, the nodes in mode $B$ contribute more to the total as each one (of the six) is involved in exactly one four-cycle (and hence raising to the power of $\alpha$ still contributes one to the sum), while of the four nodes in mode $A$, three are involved in only one four-cycle, while the fourth is involved in three four-cycles and hence contributes only $3^{\alpha}\approx 1.73205$ (when $\alpha=0.5$). Generalising this four-fan-3 graph to four-fan-$k$ ($k\geq 1$, and if $k=1$ the graph is just a four-cycle) with the central high-degree node in node set $A$, we have:
@@
Because $0<\alpha\leq 1$, the BipartiteFourCyclesNodePowerA statistic (22) will always be less than (or equal to, if $\alpha=1$ or $k=1$) the BipartiteFourCyclesNodePowerB statistic (23) for this family of graphs.
6.1 Interpretation of the new parameters
Interpretation of the BipartiteFourCyclesNodePowerA and BipartiteFourCyclesNodePowerB parameters is rather more complicated and is illustrated in Fig. 4 and Fig. 5. These figures show statistics (Fig. 4) and network visualizations (Fig. 5) of simulated bipartite networks with different combinations of positive, zero and negative BipartiteFourCyclesNodePowerA and BipartiteFourCyclesNodePowerB parameters. The simulated networks have 100 nodes in node set $A$ and 50 nodes in node set $B$ and are simulated with common ERGM parameters Edge, BipartiteAltStarsA [$\lambda=2$], and BipartiteAltStarsB [$\lambda=2$] set to $-6.0$, $-0.4$, and $1.0$, respectively. For BipartiteFourCyclesNodePowerA [$\alpha=1/5$] and BipartiteFourCyclesNodePowerB [$\alpha=1/5$] the negative (“neg”) parameter value is $-1.5$ and the positive (“pos”) parameter value is $6.5$.
Note that in a bipartite network, any four-cycle must contain two nodes in node set $A$ and two nodes in node set $B$. So how can we get more four-cycles in one mode than the other? The answer is that the four-cycle counts for the two must be equal, but the weighted node-oriented four-cycle counts (14) and (15) can differ. As discussed in Section 6, the statistic BipartiteFourCyclesNodePowerA (14) is maximised by having four-cycles involving distinct pairs of nodes in node set $A$ (rather than many four-cycles involving the same node in node set $A$). If the BipartiteFourCyclesNodePower parameter for $A$ is positive and for $B$ is zero (or negative) then we tend to get more mode $A$ nodes involved in four-cycles, with the same mode $B$ nodes participating in many four-cycles, since the statistic is higher by having different nodes in the four-cycles, than for having the same node involved in many four-cycles. In the examples illustrated in Fig. 5, this results in the mode $A$ nodes being part of a denser core with lots of four-cycles with a smaller number of $B$ nodes, resulting in isolated $B$ nodes. And vice versa for $A$ zero (or negative) and $B$ positive (“zero.pos” and “neg.pos”; these are perhaps clearer as there are more $A$ nodes than $B$ nodes in the network). Particularly in the “neg.pos” case, we can see a core of mode $B$ nodes connected to a small number of central mode $A$ nodes in four-cycles (as well as others not involved in four-cycles) and many mode $A$ isolates. If BipartiteFourCyclesNodePower for both $A$ and $B$ are positive then there are even more four-cycles, but they are more evenly distributed between the $A$ and $B$ nodes.
Figure 4: Effect of different combinations of negative, zero, and positive values of the BipartiteFourCyclesNodePowerA and BipartiteFourCyclesNodePowerB parameters on (a) the BipartiteFourCyclesNodePowerA statistic, (b) the BipartiteFourCyclesNodePowerB statistic and (c) the FourCycles statistic. Box plots show the statistics of 100 simulated networks.
To try to make this interpretation clearer, consider Fig. 6. The box plots in this figure show the number of unique nodes in each node set ($A$ or $B$) that are involved in four-cycles. When the BipartiteFourCyclesNodePowerA parameter is positive and the BipartiteFourCyclesNodePowerB parameter is zero or negative, then a large number of node set $A$ nodes are involved in four-cycles, but only a small number of node set $B$ nodes are. So the same node set $B$ nodes are involved in multiple four-cycles with many different node set $A$ nodes. If BipartiteFourCyclesNodePowerB is positive and BipartiteFourCyclesNodePowerA is negative or zero, then the same nodes in mode set $A$ are involved in multiple four-cycles with many different nodes from node set $B$.
Figure 6: For each combination of negative, zero, and positive values of the BipartiteFourCyclesNodePowerA and BipartiteFourCyclesNodePowerB parameters, there are two box plots. They show the number of unique nodes in node set A (left, red) and node set B (right, blue) involved in four-cycles. The simulations are the same as those used in Fig. 4, and as in that figure, combinations of only negative and zero values are not shown as they result in extremely low density graphs with no four-cycles (similar to the zero.zero case).
@@ 6.2 Change statistics for the new statistics
$=\sum_{k\in N(j)}L_{2}(i,k).$ (25)
The change statistic for the FourCyclesNodePower statistic (13) is then:
@@ 6.3 Position of the new statistics in the dependence hierarchy
The configurations allowed in a model are determined by the assumptions as to which ties are allowed to depend on which other ties. Pattison & Snijders [25] (subsequently elucidated by Wang et al. [22] for bipartite networks, and more recently by Pattison et al. [80]) created a two-dimensional hierarchy of dependence assumptions, where the two dimensions are two facets of proximity: the form of the proximity condition, and the maximum distance between dependent ties. This two-dimensional hierarchy of dependence assumptions is illustrated in Fig. 7, showing a partial order structure, in which, if one dependence condition can be implied by another, it can be reached by a downwards path from the first to the second [22, p. 215].
Figure 7: Dependence hierarchy, adapted from Wang et al. [22, Fig. 3]. $\mathrm{SI}_{0}$ is not defined.
In order to describe the proximity conditions, it is useful to define some notation for neighbourhoods in a graph. In Section 2 we defined $N(i)$ as the neighbours of node $i$, that is, nodes $k\neq i$ such that $d(i,k)=1$. Following Pattison et al. [80], we now define $N_{q}(u)$, the $q$-neighbourhood of node $u$, as the set of nodes within geodesic distance $q$ of $u$, that is, nodes $v$ such that $d(u,v)\leq q$. Note that $N(u)$ as earlier defined in Section 2 is distinct from $N_{1}(u)$, the former being defined as nodes with a geodesic distance of exactly 1 from $u$ (and hence excluding $u$ itself), while the latter includes $u$ itself as $d(u,u)=0$. Further, we define $N_{q}(U)$, the $q$-neighbourhood of node set $U$, as $N_{q}(U)=\{v:v\in N_{q}(u)\text{ for some }u\in U\}$, that is, the set of nodes whose distance to some node in $U$ is no more than $q$.
@@
The dependence condition $\mathrm{I}_{1}$ is equivalent to the widely-used “social circuit“ dependence assumption [22], in which two ties are conditionally dependent if they form a four-cycle if both present [81, 82, 33, 40], and $\mathrm{PI}_{1}$ allows the “alternating pendant-triangle” statistics recently described by Pattison et al. [80].
The simple but unsuccessful new statistics described in Section 5 are in the same dependence class ($\mathrm{I}_{1}$, “social circuit”) as the original $\mathrm{K\text{-}C_{A}}$ and $\mathrm{K\text{-}C_{P}}$ statistics from which they are derived.
@@
Proposition 3(d) of Pattison et al. [80] states that “For $\mathrm{D}_{\mathrm{p}}$, each configuration is a subgraph in which every pair of edges lies on a path of length $\leq(p+2)$[80, p. 191]. So for a configuration to be in $\mathrm{D}_{1}$, every pair of edges must lie on a path of length three (or shorter). The FourCyclesNodePower configuration (Fig. 3(a)) does not meet this requirement, since there are edges that do not lie on a path of length three or less: consider, for example, a pair of edges incident to the outermost node in the figure on two different four-cycles. These do not lie on a path of length three (but are on a path of length four, satisfying the requirement for $\mathrm{D}_{2}$ but not $\mathrm{D}_{1}$). Hence the FourCyclesNodePower configuration is not in dependence class $\mathrm{D}_{1}$.
Proposition 3(c) of Pattison et al. [80] states that “For $\mathrm{PI}_{\mathrm{p}}$, each configuration is a subgraph in which every pair of edges lies either on a cyclic walk of length $\leq(2p+2)$ or on a cyclic walk of length $\leq 2(p-r)+1$ with an additional path of length $\leq r+1$ attached to a node lying on the cyclic walk, for $10\leq r\leq p-$ 1[80, p. 191]. So, for $\mathrm{PI}_{2}$, each configuration is a subgraph in which every pair of edges is on a cyclic walk of length $\leq(2p+2)=6$, or on a cyclic walk of length $\leq 2(p-0)+1=5$ with an additional path of length $\leq 1$ attached to a node on the cyclic walk, or on a cyclic walk of length $\leq 2(p-1)+1=3$ with an additional path of length $\leq 2$ attached to a node on the cyclic walk. Again, we can see that the configuration for FourCyclesNodePower does not meet these conditions, considering a pair of maximally distant edges (those incident to the outermost nodes in two different four-cycles in Fig. 3(a)). Such a pair of edges is neither on a six-cycle, and nor is it on a five-cycle with an additional path of length at most one or a three-cycle with an additional path of length at most two. Hence the FourCyclesNodePower configuration is not in dependence class $\mathrm{PI}_{2}$.
In the case of the BipartiteFourCyclesNodePower ($A$ and $B$) statistics for bipartite networks, the same reasoning applies (see Fig. 3(b) and Fig. 3(c)).
@@ 6.4 Implementation
The new ERGM effects FourCyclesNodePower, BipartiteFourCyclesNodePowerA, and BipartiteFourCyclesNodePowerB are implemented in the EstimNetDirected [74] software, available from https://github.com/stivalaa/EstimNetDirected. BipartiteFourCyclesNodePowerA and BipartiteFourCyclesNodePowerB are also implemented, as b1np4c and b2np4c, as user-contributed statnet model terms [83, 84], available from https://github.com/stivalaa/ergm.terms.contrib. An example is described in Section 8.
@@ 7 Simulation experiments
The results of these simulations are shown in Fig. 8. Using the BipartiteAltKCyclesA parameter (left column) results in phase transition or near-degeneracy behaviour, with the statistic showing a sudden sharp increase at a critical value of the parameter. At this critical value, the graph density (Fig. 8(a) left plot) also sharply increases, as does the number of four-cycles (Fig. 8(c) left plot), which, at parameter values less than the critical value, hardly increased at all. This behaviour is similar to that of the simple Markov (edge-triangle) model described in Koskinen & Daraganova [40], and is characteristic of near-degeneracy in ERGMs. This can prevent estimation of models which contain parameters that cause this behaviour, and yet occurs in this case even when using an “alternating” statistic, designed to try to avoid such behaviour [21]. Note that changing the $\lambda$ parameter appears merely to change the maximum value of the statistic; it does not remove or “smooth out” the phase transition. In contrast, when using the new BipartiteFourCyclesNodePowerA parameter (right column) the phase transition behaviour is less apparent even for the highest value of $\alpha$ ($1/2$), and can be smoothed out further as we decrease $\alpha$. So, by appropriately setting $\alpha$, we can use the BipartiteFourCyclesNodePowerA parameter to generate smoothly varying numbers of four-cycles, without suddenly tipping from a low-density low-clustering regime to a high-density high-clustering regime with nothing in between, which did not seem to be possible on this example with the BipartiteAltKCyclesA parameter.